Sir/Madam This is the format of 4 numbers in AP as (a-3d),(a-d),(a+d),(a+3d) is given in the book. I want to know “How this AP formed”.Please explain this concept . Thanks & Regards Monika

Sir/Madam
This  is the format of 4 numbers in AP as (a-3d),(a-d),(a+d),(a+3d) is given in the book.
I want to know “How this AP formed”.Please explain this concept .
Thanks & Regards
Monika

Grade:

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
here
a1= a-3d
a2= a-d
a3=a+d
a4=a+3d

Find the difference between two consecutive terms
a2-a1=2d
a3-a2=2d
a4-a3=2d

Since difference is constant and I guess you know the definition of in AP.